OpenStudy (peterg2001):

Choose the answer based on the most efficient method as presented in the lesson. If the first step in the solution of the equation -8 - 7x = -5x - 10 is "add 10," then what should the next step be? add 5x add 7x add 8

OpenStudy (peterg2001):

@haleyelizabeth2017

Directrix (directrix):

-8 - 7x = -5x - 10 If you add 10 to both sides, you get: -8 + 10 - 7x = -5x - 10 + 10 which is 2 - 7x = -5x The idea in solving linear equations is to collect the variables on one side of the equation and the constant terms on the other and then divide. In light of that and looking at the options, what would you next add to both sides of the equation? @peterg2001 2 - 7x = -5x